# Calibrating Long-form Generations from Large Language Models

### v1

@@ Appendix E Prompts
E.1 Correctness Evaluation Template
Similar to contemporary work (Kim et al., 2023), our evaluation template for GPT-4 evaluation to get the ground-truth correctness distribution of an answer includes four components: a clear task description, expertly crafted evaluation criteria for objectivity, demonstrations with a variety of answer qualities (best, worst, intermediate) each with a score and rationale, and specific evaluation instructions for the LLM, encompassing the question-answer pair to be evaluated and a reference answer.
{task instruction} You will be given a question, a reference answer, and a student’s answer. Please evaluate the student’s answer based on both your knowledge and the reference answer, and provide a score from 0-5 to the student’s answer. Keep in mind that the reference answer is not the sole correct response. Assess for both factual accuracy and relevance to the question. The following are the scoring criterion: {criterion} Here are some examples. {examples} Now it’s your turn. Question: {question} Reference answer: {reference answer} Student’s answer: {answer} Now please provide your score about this answer in the format of “Score: <Your score>/5” and give your explanation.
E.2 Criteria
@@
5 - Completely Correct and Highly Relevant: The answer fully addresses the question, resolves the ambiguity, and provides a well-rounded resolution. All facts presented in the answer are accurate and relevant. 4 - Mostly Correct and Relevant: The answer is very relevant and addresses the ambiguity well, but might have a minor oversight or inaccuracy. All the facts presented are accurate and relevant, or with only minor errors. 3 - Partially Correct and Relevant: The answer is generally on topic and attempts to address the ambiguity, but there might be inaccuracies or omissions. The majority of the facts are correct, with a few errors. 2 - Flawed but Somewhat Relevant: The answer somewhat addresses the topic but does not fully explore the question’s ambiguity or does not provide a complete resolution. The facts presented are a mix of correct and incorrect information, with about half being accurate. 1 - Mostly Incorrect or Mostly Irrelevant: The answer slightly touches upon the topic but misses the main point. The majority of the facts presented are incorrect, with only a small portion being accurate. 0 - Completely Incorrect or Completely Irrelevant: The student’s answer is completely off-topic, not related to the question at all, or contains only incorrect information.
@@
5 - Perfectly Addressed, Accurate and Clarity: The answer flawlessly addresses the question with exceptional accuracy and clarity. It simplifies complex concepts effectively and does so in a way that is captivating and memorable. 4 - Accurate and clear: The answer is accurate, relevant to the question, and presented in a way that is engaging and understandable. It simplifies complex concepts effectively but may miss a small opportunity for further clarification or engagement. 3 - Moderately Accurate and Understandable: The answer is mostly accurate and somewhat understandable. It addresses the question reasonably well but may lack detail or contain some inaccuracies. It may use complex terms or concepts that are not broken down into simpler ideas. 2 - Relevant but Lacks Clarity or Accuracy: The answer is related to the question but lacks clarity or contains partial inaccuracies. It attempts to simplify the idea but does not do so effectively, leaving room for confusion or misunderstanding. 1 - Significantly Flawed: The answer addresses the question to a minimal extent but contains significant inaccuracies or misleading information. It might show a basic attempt to simplify the concept but fail in accuracy or relevance. 0 - Completely Inaccurate or Irrelevant: The answer is entirely off-topic, irrelevant, or factually incorrect. It fails to address the question and does not simplify complex ideas.
@@
* Estimated Accuracy Assessment (0-3 Points) * 3 Points: All answers provided seem correct based on available knowledge. * 2 Points: Majority of the answers seem correct. * 1 Point: Some answers are correct, but there are notable inaccuracies. * 0 Points: No answers seem correct or very high degree of inaccuracy. * Estimated Completeness Assessment (0-2 Points) * 2 Points: The response seems comprehensive, covering a broad range of known or expected correct answers. * 1 Point: The response covers some correct answers but misses significant known or expected answers. * 0 Points: The response is highly incomplete, missing most of the known or expected correct answers. * Total Score (0-5 Scale) * Add the points from Estimated Accuracy and Estimated Completeness.
@@
You will be given one summary written for a news article. Your task is to rate the overall quality of the summary with a score from 0 to 5, where 0 is the lowest and 5 is the highest. Please make sure you read and understand these instructions carefully. Please keep this document open while reviewing, and refer to it as needed. Evaluation Steps: 1. Read the news article carefully and identify the main topic and key points. 2. Read the summary and compare it to the news article. Check if the summary covers the main topic and key points of the news article, and if it presents them in a clear and logical order. 3. Assign a score for the summary quality on a scale of 0 to 5, where 0 is the lowest and 5 is the highest. Task Input: Article: {article} Summary: {summary} Now please provide your score of the summary in the format of “Score: <Your score>/5” and give your explanation.
E.3 Self-Evaluation Template
@@
{task instruction} You will be given a question and a student’s answer. Please evaluate the student’s answer, and provide a score from 0-5 to the student’s answer based on the following scoring criterion: {criterion} Here are some examples. {examples} Now it’s your turn. Question: {question} Answer: {answer} Now please provide your score about this answer in the format of “Score: <Your score>/5” and give your explanation. Assess for both factual accuracy and relevance to the question.
E.4 Self-Consistency Template
@@
You will be presented with a question followed by two answers. Evaluate how similar these answers are, considering their amount of information provided, factual content, effectiveness in addressing the question, format and organization. Conclude by providing an integer similarity score between 0 and 5. Question: {question} Answer 1: {answer1} Answer 2: {answer2} Now please give your similarity score in the format of “Similarity score: <Your score>/5” and give your explanation. Make your answer short and concise.
You will receive a sentence and a response; please ascertain if a similar statement is present in the response. Sentence: {sentence} Response: {response} Verify if the response contains a statement resembling the target sentence. Answer yes or no.
Appendix F Qualitative Analysis
We showcase some successful and failed calibration examples from LLama-2-13b on Qampari and ASQA with self-consistency method.
@@ Appendix F Qualitative Analysis
Question
Who directed a film that had P. Balachandran as a screenwriter?
Answer
Kamal, P. Balachandran, T. K. Rajeev Kumar, V. K. Prakash, Rajeev Ravi.
Model Answer
Adu, Ammayum Muyalum, Bhoothakkannadi, Chattambi, Kakkakku Pinnaka…
Other Answer 1
Agraharathile Kazhuthai, Kaliyugam, Kodungallur, Kodiyettam …
Other Answer 2
Kodiyettam, Kuttyedam, Oru CBI Diary Kurippu…
Other Answer 3
Aravindan, K.G. George, John Abraham, Shyam Benegal …
Self-consistency scores
[0, 0, 0, …] (%)
Correctness
Score: $0\%$ Distribution: $\{0:1.0\}$
Confidence score
Score: $0\%$ Distribution: $\{0:1.0\}$
@@
Question
Which cities or municipalities are located in the province of Basilan?
Answer
Akbar, Al-Barka, Hadji Muhtamad, Isabela City, Lamitan, Lantawan, Maluso, Sumisip, Tabuan-Lasa, Tipo-Tipo, Tuburan, Ungkaya Pukan.
Model Answer
Al-Barka, Basilan, Isabela, Lamitan, Tipo-Tipo, Maluso, Sumisip.
Other Answer 1
Basilan, Isabela, Lamitan, Lantawan, Tipo-Tipo, Sumisip.
Other Answer 2
Basilan, Tipo-Tipo, Isabela City, Maluso, Sumisip, Tuburan.
Other Answer 3
Basilan, Isabela, Tipo-Tipo, Lamitan, Sumisip, Al-Barka, Maluso, Tuburan.
Self-consistency scores
[71.4, 57.1, 85.7, …](%)
Correctness
Score: $83.3\%$ Distribution: $\{0.8:1.0\}$
Confidence
Score: $76.2\%$ Distribution: $\{0.6:0.11,0.8:0.78,1.0:0.11\}$
@@
Question
What city or municipality located in the Philipines are in the Tarlac province?
Answer
Anao, Bamban, Camiling, Capas, Concepcion, Gerona, La Paz, Mayantoc, Moncada, Paniqui, Pura, Ramos, San Clemente, San Jose, San Manuel, Santa Ignacia, Tarlac City, Victoria.
Model Answer
Capas, Concepcion, Gerona, La Paz, Mayantoc, Paniqui, San Jose, Santa Ignacia, Tarlac City, Victoria.
Other Answer 1
Capas, Concepcion, Camiling, Tarlac City, Paniqui, San Jose, Santa Lucia, Victoria.
Other Answer 2
Capas, Tarlac, Concepcion, Tarlac, Gerona, Tarlac, La Paz, Tarlac, Paniqui, Tarlac, San Jose, Tarlac, Tarlac City.
Other Answer 3
Capas, Concepcion, Camiling, Tarlac City, San Jose.
Self-consistency scores
[60, 60, 30, …](%)
Correctness
Score: $100\%$ Distribution: $\{1:1.0\}$
Confidence
Score: $51.1\%$ Distribution: $\{0.4:0.56,0.6:0.33,0.8:0.11\}$
@@
Question
What are some notables mailing list softwares ?
Answer
Dada Mail, Discourse, GNU Mailman, GroupServer, LISTSERV, Majordomo, Organic Groups Mailing List, Sympa.
Model Answer
Mailchimp, Constant Contact, Campaign Monitor, AWeber, GetResponse, MailPoet, Sendinblue, MailerLite, Klaviyo, ConvertKit.
Other Answer 1
Mailchimp, AWeber, GetResponse, Campaign Monitor, Constant Contact, Klaviyo, ConvertKit, MailPoet, Mailster, Sendinblue, MailerLite.
Other Answer 2
Mailchimp, Constant Contact, AWeber, Campaign Monitor, GetResponse, Klaviyo, MailPoet, Mailster, Sendinblue, MailerLite, ConvertKit, ActiveCampaign, and Constant Contact.
Other Answer 3
Mailchimp, AWeber, GetResponse, Campaign Monitor, MailerLite, Constant Contact, Klaviyo, ConvertKit, Sendinblue, and Emma.
Self-consistency scores
[100, 100, 90, …](%)
Correctness
Score: $0\%$ Distribution: $\{0:1.0\}$
Confidence score
Score: $97.8\%$ Distribution: $\{0.8:0.22,1.0:0.78\}$
@@
Question
What are the three types of enzymes in digestion?
Answer
Digestive enzymes are a group of enzymes that break down polymeric… Lipases split fatty acids… Proteases and peptidases split proteins … Amylases split carbohydrates … Nucleases split nucleic acids into nucleotides…
Model Answer
There are three main types of enzymes involved in digestion: proteases, which break down proteins; amylases, which break down carbohydrates; and lipases, which break down fats. Proteases are produced by the stomach and pancreas, while amylases are … These enzymes work in concert …
Other Answer 1
Digestive enzymes are proteins that help … There are three main types of digestive enzymes: proteases, which break down proteins into amino acids; amylases, which break down carbohydrates into simple sugars; and lipases, which break down fats into glycerol and fatty acids. These enzymes are produced by different organs and …
Other Answer 2
There are three main types of enzymes involved in digestion: proteases, which break down proteins; amylases, which break down carbohydrates; and lipases, which break down fats. These enzymes are produced by the salivary glands, stomach, and pancreas, and work in various parts of the digestive system…
Self-consistency scores
[100, 100, 0, …] (%)
Correctness
Score: $100\%$ Distribution: $\{1.0:1.0\}$
Confidence score
Score: $95.6\%$ Distribution: $\{0.8:0.22,1.0:0.78\}$
@@
Question
When did ben rectors new album come out?
Answer
Benjamin Evans Rector is an American singer… He has released seven studio albums, including Magic on June 22, 2018, which peaked at No. 1 .. Brand New which released on August 28, 2015… Older albums such as Something Like This, released on September 13, 2011 and The Walking in Between, came out August 20, 2013.
Model Answer
Ben Rector’s new album, "Magic", was released on March 25, 2022.
Other Answer 1
Ben Rector’s new album, "Brand New", was released on March 10th, 2022.
Other Answer 2
Ben Rector, an American singer-songwriter, released his new album "Magic" on October 20, 2022.
Self-consistency scores
[0, 0, …](%)
Correctness
Score: $0\%$ Distribution: $\{0:1.0\}$
Confidence
Score: $0\%$ Distribution: $\{0:1.0\}$
@@
Question
Who plays charlie on it’s always sunny?
Answer
Charlie Day plays the fictional character Charlie Kelly… Day is an American actor, screenwriter…
Model Answer
Charlie Day plays Charlie on It’s Always Sunny in Philadelphia. He is also known for his role as Dale Arbus … Day is a member of the improvisational comedy …
Other Answer 1
Dennis is played by Glenn Howerton.
Other Answer 2
Charlie on It’s Always Sunny in Philadelphia is played by Charlie Day.
Self-consistency scores
[0, 33.3, …](%)
Correctness
Score: $80\%$ Distribution: $\{0.8:1.0\}$
Confidence
Score: $31.1\%$ Distribution: $\{0:0.22,0.4:0.78\}$
@@
Question
Who sings wake me up when it’s all over?
Answer
“Wake Me Up” is a song by Swedish DJ and record producer Avicii. Aloe Blacc, which is the stage name for Egbert Nathaniel Dawkins III provides vocals for the song while Mike Einziger provides acoustic guitar. “Wake Me When It’s Over” is a song written and recorded by country singer Willie Nelson in 1962.
Model Answer
“Wake Me Up When September Ends” is a song written by Billy
Other Answer 1
“Wake Me Up When September Ends” is a song by American rock band Green Day, written by lead vocalist Billie Joe Armstrong…
Other Answer 2
“Wake Me Up When September Ends” is a song by Green Day, written by the band’s lead vocalist and guitarist Billie Joe Armstrong…
Self-consistency scores
[100, 100, …](%)
Correctness
Score: $0\%$ Distribution: $\{0:1.0\}$
Confidence score
Score: $77.8\%$ Distribution: $\{0:0.22,1.0:0.78\}$

### v2

@@ Appendix F Prompts
F.1 Correctness Evaluation Template
Similar to contemporary work (Kim et al., 2023), our evaluation template for GPT-4 evaluation to get the target correctness distribution of an answer includes four components: a clear task description, expertly crafted evaluation criteria for objectivity, demonstrations with a variety of answer qualities (best, worst, intermediate) each with a score and rationale, and specific evaluation instructions for the LLM, encompassing the question-answer pair to be evaluated and a reference answer.
Evaluation Template {task instruction} You will be given a question, a reference answer, and a student’s answer. Please evaluate the student’s answer based on both your knowledge and the reference answer, and provide a score from 0-5 to the student’s answer. Keep in mind that the reference answer is not the sole correct response. Assess for both factual accuracy and relevance to the question. The following are the scoring criterion: {criterion} Here are some examples. {examples} Now it’s your turn. Question: {question} Reference answer: {reference answer} Student’s answer: {answer} Now please provide your score about this answer in the format of “Score: <Your score>/5” and give your explanation.
F.2 Criteria
@@
ASQA Criterion 5 - Completely Correct and Highly Relevant: The answer fully addresses the question, resolves the ambiguity, and provides a well-rounded resolution. All facts presented in the answer are accurate and relevant. 4 - Mostly Correct and Relevant: The answer is very relevant and addresses the ambiguity well, but might have a minor oversight or inaccuracy. All the facts presented are accurate and relevant, or with only minor errors. 3 - Partially Correct and Relevant: The answer is generally on topic and attempts to address the ambiguity, but there might be inaccuracies or omissions. The majority of the facts are correct, with a few errors. 2 - Flawed but Somewhat Relevant: The answer somewhat addresses the topic but does not fully explore the question’s ambiguity or does not provide a complete resolution. The facts presented are a mix of correct and incorrect information, with about half being accurate. 1 - Mostly Incorrect or Mostly Irrelevant: The answer slightly touches upon the topic but misses the main point. The majority of the facts presented are incorrect, with only a small portion being accurate. 0 - Completely Incorrect or Completely Irrelevant: The student’s answer is completely off-topic, not related to the question at all, or contains only incorrect information.
@@
ELI5 Criterion 5 - Perfectly Addressed, Accurate and Clarity: The answer flawlessly addresses the question with exceptional accuracy and clarity. It simplifies complex concepts effectively and does so in a way that is captivating and memorable. 4 - Accurate and clear: The answer is accurate, relevant to the question, and presented in a way that is engaging and understandable. It simplifies complex concepts effectively but may miss a small opportunity for further clarification or engagement. 3 - Moderately Accurate and Understandable: The answer is mostly accurate and somewhat understandable. It addresses the question reasonably well but may lack detail or contain some inaccuracies. It may use complex terms or concepts that are not broken down into simpler ideas. 2 - Relevant but Lacks Clarity or Accuracy: The answer is related to the question but lacks clarity or contains partial inaccuracies. It attempts to simplify the idea but does not do so effectively, leaving room for confusion or misunderstanding. 1 - Significantly Flawed: The answer addresses the question to a minimal extent but contains significant inaccuracies or misleading information. It might show a basic attempt to simplify the concept but fail in accuracy or relevance. 0 - Completely Inaccurate or Irrelevant: The answer is entirely off-topic, irrelevant, or factually incorrect. It fails to address the question and does not simplify complex ideas.
@@
QAMPARI Criterion * Estimated Accuracy Assessment (0-3 Points) * 3 Points: All answers provided seem correct based on available knowledge. * 2 Points: Majority of the answers seem correct. * 1 Point: Some answers are correct, but there are notable inaccuracies. * 0 Points: No answers seem correct or very high degree of inaccuracy. * Estimated Completeness Assessment (0-2 Points) * 2 Points: The response seems comprehensive, covering a broad range of known or expected correct answers. * 1 Point: The response covers some correct answers but misses significant known or expected answers. * 0 Points: The response is highly incomplete, missing most of the known or expected correct answers. * Total Score (0-5 Scale) * Add the points from Estimated Accuracy and Estimated Completeness.
@@
Summarization Evaluation Template You will be given one summary written for a news article. Your task is to rate the overall quality of the summary with a score from 0 to 5, where 0 is the lowest and 5 is the highest. Please make sure you read and understand these instructions carefully. Please keep this document open while reviewing, and refer to it as needed. Evaluation Steps: 1. Read the news article carefully and identify the main topic and key points. 2. Read the summary and compare it to the news article. Check if the summary covers the main topic and key points of the news article, and if it presents them in a clear and logical order. 3. Assign a score for the summary quality on a scale of 0 to 5, where 0 is the lowest and 5 is the highest. Task Input: Article: {article} Summary: {summary} Now please provide your score of the summary in the format of “Score: <Your score>/5” and give your explanation.
F.3 Self-Evaluation Template
@@
Self Evaluation Template {task instruction} You will be given a question and a student’s answer. Please evaluate the student’s answer, and provide a score from 0-5 to the student’s answer based on the following scoring criterion: {criterion} Here are some examples. {examples} Now it’s your turn. Question: {question} Answer: {answer} Now please provide your score about this answer in the format of “Score: <Your score>/5” and give your explanation. Assess for both factual accuracy and relevance to the question.
F.4 Self-Consistency Template
@@
Similarity Comparison Naive Template You will be presented with a question followed by two answers. Evaluate how similar these answers are, considering their amount of information provided, factual content, effectiveness in addressing the question, format and organization. Conclude by providing an integer similarity score between 0 and 5. Question: {question} Answer 1: {answer1} Answer 2: {answer2} Now please give your similarity score in the format of “Similarity score: <Your score>/5” and give your explanation. Make your answer short and concise.
Similarity Comparison Split Template You will receive a sentence and a response; please ascertain if a similar statement is present in the response. Sentence: {sentence} Response: {response} Verify if the response contains a statement resembling the target sentence. Answer yes or no.
Appendix G Qualitative Analysis
We showcase successful and failed calibration examples on Qampari and ASQA with self-consistency method.
@@ Appendix G Qualitative Analysis
Question Who directed a film that had P. Balachandran as a screenwriter?
Answer Kamal, P. Balachandran, T. K. Rajeev Kumar, V. K. Prakash, Rajeev Ravi.
Model Answer Adu, Ammayum Muyalum, Bhoothakkannadi, Chattambi, Kakkakku Pinnaka…
Other Answer 1 Agraharathile Kazhuthai, Kaliyugam, Kodungallur, Kodiyettam …
Other Answer 2 Kodiyettam, Kuttyedam, Oru CBI Diary Kurippu…
Other Answer 3 Aravindan, K.G. George, John Abraham, Shyam Benegal …
Self-consistency scores [0, 0, 0, …] (%)
Correctness Score: $0\%$ Distribution: $\{0:1.0\}$
Confidence score Score: $0\%$ Distribution: $\{0:1.0\}$
@@
Question Which cities or municipalities are located in the province of Basilan?
Answer Akbar, Al-Barka, Hadji Muhtamad, Isabela City, Lamitan, Lantawan, Maluso, Sumisip, Tabuan-Lasa, Tipo-Tipo, Tuburan, Ungkaya Pukan.
Model Answer Al-Barka, Basilan, Isabela, Lamitan, Tipo-Tipo, Maluso, Sumisip.
Other Answer 1 Basilan, Isabela, Lamitan, Lantawan, Tipo-Tipo, Sumisip.
Other Answer 2 Basilan, Tipo-Tipo, Isabela City, Maluso, Sumisip, Tuburan.
Other Answer 3 Basilan, Isabela, Tipo-Tipo, Lamitan, Sumisip, Al-Barka, Maluso, Tuburan.
Self-consistency scores [71.4, 57.1, 85.7, …](%)
Correctness Score: $83.3\%$ Distribution: $\{0.8:1.0\}$
Confidence Score: $76.2\%$ Distribution: $\{0.6:0.11,0.8:0.78,1.0:0.11\}$
@@
Question What city or municipality located in the Philipines are in the Tarlac province?
Answer Anao, Bamban, Camiling, Capas, Concepcion, Gerona, La Paz, Mayantoc, Moncada, Paniqui, Pura, Ramos, San Clemente, San Jose, San Manuel, Santa Ignacia, Tarlac City, Victoria.
Model Answer Capas, Concepcion, Gerona, La Paz, Mayantoc, Paniqui, San Jose, Santa Ignacia, Tarlac City, Victoria.
Other Answer 1 Capas, Concepcion, Camiling, Tarlac City, Paniqui, San Jose, Santa Lucia, Victoria.
Other Answer 2 Capas, Tarlac, Concepcion, Tarlac, Gerona, Tarlac, La Paz, Tarlac, Paniqui, Tarlac, San Jose, Tarlac, Tarlac City.
Other Answer 3 Capas, Concepcion, Camiling, Tarlac City, San Jose.
Self-consistency scores [60, 60, 30, …](%)
Correctness Score: $100\%$ Distribution: $\{1:1.0\}$
Confidence Score: $51.1\%$ Distribution: $\{0.4:0.56,0.6:0.33,0.8:0.11\}$
@@
Question What are some notables mailing list softwares ?
Answer Dada Mail, Discourse, GNU Mailman, GroupServer, LISTSERV, Majordomo, Organic Groups Mailing List, Sympa.
Model Answer Mailchimp, Constant Contact, Campaign Monitor, AWeber, GetResponse, MailPoet, Sendinblue, MailerLite, Klaviyo, ConvertKit.
Other Answer 1 Mailchimp, AWeber, GetResponse, Campaign Monitor, Constant Contact, Klaviyo, ConvertKit, MailPoet, Mailster, Sendinblue, MailerLite.
Other Answer 2 Mailchimp, Constant Contact, AWeber, Campaign Monitor, GetResponse, Klaviyo, MailPoet, Mailster, Sendinblue, MailerLite, ConvertKit, ActiveCampaign, and Constant Contact.
Other Answer 3 Mailchimp, AWeber, GetResponse, Campaign Monitor, MailerLite, Constant Contact, Klaviyo, ConvertKit, Sendinblue, and Emma.
Self-consistency scores [100, 100, 90, …](%)
Correctness Score: $0\%$ Distribution: $\{0:1.0\}$
Confidence score Score: $97.8\%$ Distribution: $\{0.8:0.22,1.0:0.78\}$
@@
Question What are the three types of enzymes in digestion?
Answer Digestive enzymes are a group of enzymes that break down polymeric… Lipases split fatty acids… Proteases and peptidases split proteins … Amylases split carbohydrates … Nucleases split nucleic acids into nucleotides…
Model Answer There are three main types of enzymes involved in digestion: proteases, which break down proteins; amylases, which break down carbohydrates; and lipases, which break down fats. Proteases are produced by the stomach and pancreas, while amylases are … These enzymes work in concert …
Other Answer 1 Digestive enzymes are proteins that help … There are three main types of digestive enzymes: proteases, which break down proteins into amino acids; amylases, which break down carbohydrates into simple sugars; and lipases, which break down fats into glycerol and fatty acids. These enzymes are produced by different organs and …
Other Answer 2 There are three main types of enzymes involved in digestion: proteases, which break down proteins; amylases, which break down carbohydrates; and lipases, which break down fats. These enzymes are produced by the salivary glands, stomach, and pancreas, and work in various parts of the digestive system…
Self-consistency scores [100, 100, 0, …] (%)
Correctness Score: $100\%$ Distribution: $\{1.0:1.0\}$
Confidence score Score: $95.6\%$ Distribution: $\{0.8:0.22,1.0:0.78\}$
@@
Question When did ben rectors new album come out?
Answer Benjamin Evans Rector is an American singer… He has released seven studio albums, including Magic on June 22, 2018, which peaked at No. 1 .. Brand New which released on August 28, 2015… Older albums such as Something Like This, released on September 13, 2011 and The Walking in Between, came out August 20, 2013.
Model Answer Ben Rector’s new album, "Magic", was released on March 25, 2022.
Other Answer 1 Ben Rector’s new album, "Brand New", was released on March 10th, 2022.
Other Answer 2 Ben Rector, an American singer-songwriter, released his new album "Magic" on October 20…
Self-consistency scores [0, 0, …](%)
Correctness Score: $0\%$ Distribution: $\{0:1.0\}$
Confidence Score: $0\%$ Distribution: $\{0:1.0\}$
@@
Question Who plays charlie on it’s always sunny?
Answer Charlie Day plays the fictional character Charlie Kelly… Day is an American actor, screenwriter…
Model Answer Charlie Day plays Charlie on It’s Always Sunny in Philadelphia. He is also known for his role as Dale Arbus … Day is a member of the improvisational comedy …
Other Answer 1 Dennis is played by Glenn Howerton.
Other Answer 2 Charlie on It’s Always Sunny in Philadelphia is played by Charlie Day.
Self-consistency scores [0, 33.3, …](%)
Correctness Score: $80\%$ Distribution: $\{0.8:1.0\}$
Confidence Score: $31.1\%$ Distribution: $\{0:0.22,0.4:0.78\}$
@@
Question Who sings wake me up when it’s all over?
Answer “Wake Me Up” is a song by Swedish DJ and record producer Avicii. Aloe Blacc, which is the stage name for Egbert Nathaniel Dawkins III provides vocals for the song while Mike Einziger provides acoustic guitar. “Wake Me When It’s Over” is a song written and recorded by Willie…
Model Answer “Wake Me Up When September Ends” is a song written by Billy
Other Answer 1 “Wake Me Up When September Ends” is a song by American rock band Green Day, written by lead vocalist Billie Joe Armstrong…
Other Answer 2 “Wake Me Up When September Ends” is a song by Green Day, written by the band’s lead vocalist and guitarist Billie Joe Armstrong…
Self-consistency scores [100, 100, …](%)
Correctness Score: $0\%$ Distribution: $\{0:1.0\}$
Confidence score Score: $77.8\%$ Distribution: $\{0:0.22,1.0:0.78\}$